OpenStudy (anonymous):

Which Property of Equality should be used first to solve the equation p divided by seven minus five equals negative three for p Addition Property of Equality Subtraction Property of Equality Multiplication Property of Equality Division Property of Equality

OpenStudy (anonymous):

Division and Multiplication are always used before Addition and Subtraction. Since there is Division and no Multiplication... The answer is: Division Property of Equality

OpenStudy (anonymous):

Assuming that your equation is this: \[ \frac{ P }{ 7 } - 5 = -3\] *Thats how I read it

OpenStudy (anonymous):

If your equation is really this: \[\frac{ P }{ 7-5 } = -3\] Then the Subtraction Property of Equality would be used first.